OpenStudy (anonymous):

Please help? Carla paid $462.25 for textbooks with 7.5% sales tax included. What was the price of the textbooks?

OpenStudy (anonymous):

496.92

OpenStudy (anonymous):

thats not right

OpenStudy (anonymous):

the $462.25 is what she paid for the book including tax 7.5%

OpenStudy (anonymous):

Oh!

OpenStudy (anonymous):

Hold on i'll solve

OpenStudy (anonymous):

462.25=.925(p) P=Original Price. Divide both sides by .925 427.58125 =$427.58

OpenStudy (anonymous):

Translate the following situation into an equation. Do not solve. “The sum of three consecutive odd integers is 85 less than four times the first. What is the first integer?”

OpenStudy (anonymous):

Let n = first integer n + 1 = second integer n + 2 = third integer n + (n + 1) + (n + 2) = 4n - 85 3n + 3 = 4n - 85 88 = n